Evaluate the integral of Sin (5/2) x dx


1
Expert's answer
2021-03-10T12:06:19-0500

"\int sin(5/2)x dx = sin(5/2) \times \int xdx = sin(5/2) \frac{x^{1+1}}{1+1} = \frac{1}{2}sin(5/2)x^2 + C"